# On the metric of the jet bundle and similarity on weighted Dirichlet spaces

### v2

@@ 3. Similarity in the class $\mathcal{B}_{m}(\mathbb{D})$
where $E=\bigvee\limits_{w\in\mathbb{D}}\{\widetilde{\gamma}_{i}(w):1\leq i\leq m\}$. Obviously, the function $F^{\#}$ defined on $\mathbb{D}$ as $F^{\#}(w):=F(\overline{w})$ is such that $(F^{\#})^{*}\in\text{Mult}(\mathcal{M}_{n,E},\mathcal{M}_{n,\mathbb{C}^{m}}).$ By Theorem 2.1,
@@
Theorem 2.4 in [16] and Theorem 2.3 in [24] use the curvature of the holomorphic complex bundles to characterize the similarity of $n$-hypercontraction. Next, we give the following brief and concise proof using Theorem 3.4:
@@
Let $T\in\mathcal{B}_{1}(\mathbb{D})$ be an $n$-hypercontraction. Then $T$ is similar to the backward shift operator $S^{*}_{n}$ on $\mathcal{M}_{n}$ if and only if there exists a bounded subharmonic function $\varphi$ over $\mathbb{D}$ such that
$\mathcal{K}_{S^{*}_{n}}(w)-\mathcal{K}_{T}(w)=\frac{\partial^{2}}{\partial w\partial\overline{w}}\varphi(w),\quad w\in\mathbb{D}.$
@@
Since $T\in\mathcal{B}_{1}(\mathbb{D})$ is an $n$-hypercontraction, by Lemma 3.2, there exists a Hermitian holomorphic line bundle $\mathcal{E}_{D_{T}}$ over $\mathbb{D}$ such that $\mathcal{E}_{T}\sim_{u}\mathcal{E}_{S^{*}_{n}}\otimes\mathcal{E}_{D_{T}}$.
If $T$ is similar to $S^{*}_{n}$, there is a bounded invertible operator $X$ such that
$X\left(\mathcal{E}_{S^{*}_{n}}(w)\right)=\left(\mathcal{E}_{S^{*}_{n}}\otimes\mathcal{E}_{D_{T}}\right)(w),\quad w\in\mathbb{D}.$
Therefore, there exist constants $C_{1}$, $C_{2}$, and a non-vanishing holomorphic cross-section $\gamma$ of $\mathcal{E}_{D_{T}}$ satisfying
$XK_{n}(\cdot,\overline{w})=K_{n}(\cdot,\overline{w})\otimes\gamma(w)\quad\text{and}\quad 0<C_{1}\leq\|\gamma(w)\|^{2}\leq C_{2},\quad w\in\mathbb{D}.$
Hence,
$\mathcal{K}_{S^{*}_{n}}(w)-\mathcal{K}_{T}(w)=\frac{\partial^{2}}{\partial w\partial\overline{w}}\log\|\gamma(w)\|^{2}=\frac{\|\gamma(w)\|^{2}\frac{\partial^{2}}{\partial w\partial\overline{w}}\|\gamma(w)\|^{2}-\frac{\partial}{\partial w}\|\gamma(w)\|^{2}\frac{\partial}{\partial\overline{w}}\|\gamma(w)\|^{2}}{\|\gamma(w)\|^{4}}\geq 0.$
To complete the proof, one sets $\varphi(w):=\log\|\gamma(w)\|^{2}$ on $\mathbb{D}$.
For the other implication, we first note that since $\mathcal{E}_{T}\sim_{u}\mathcal{E}_{S^{*}_{n}}\otimes\mathcal{E}_{D_{T}}$ and $\mathcal{K}_{T}=\mathcal{K}_{S^{*}_{n}}+\mathcal{K}_{\mathcal{E}_{D_{T}}}.$ Then
$\frac{\partial^{2}}{\partial w\partial\overline{w}}\varphi(w)=\frac{\partial^{2}}{\partial w\partial\overline{w}}\log\|{\gamma}(w)\|^{2}$
for bounded subharmonic function $\varphi$ on $\mathbb{D}$ and a non-vanishing holomorphic cross-section ${\gamma}$ of $\mathcal{E}_{D_{T}}$. Now, $u(w):=\log\left(\frac{\|{\gamma}(w)\|^{2}}{e^{\varphi(w)}}\right)^{\frac{1}{2}}$ is a real-valued harmonic function, and letting $\phi:=e^{u+iv}\in\text{Hol}(\mathbb{D}),$ where $v$ is the harmonic conjugate of $u$, we see that
@@
Then $e^{{\varphi}(w)}=\left\|\frac{1}{\phi(w)}{\gamma}(w)\right\|^{2}$ and $\frac{1}{\phi}{\gamma}$ is a holomorphic cross-section of $\mathcal{E}_{D_{T}}$. Since $\varphi$ is bounded on $\mathbb{D}$, there are constants $C_{1}$ and $C_{2}$ such that
@@
and therefore by Theorem 3.4, $T\sim_{s}S^{*}_{n}$. ∎
Two Cowen-Douglas operators are unitarily equivalent if and only if their corresponding holomorphic complex bundles are equivalent as Hermitian holomorphic vector bundles. It is known from [40] that many holomorphic complex bundles with tensor products correspond to the Cowen-Douglas operator. Therefore, we use holomorphic complex bundles to describe the similar classification of the Cowen-Douglas operators and obtain the following proposition, whose proof is similar to Theorem 3.4, so the proof is omitted.
@@
Let $T\in\mathcal{B}_{m}(\mathbb{D})$ and $M_{z}^{*}$ be the backward shift operator on reproducing kernel Hilbert space $\mathcal{H}_{K}$ with $\text{Mult}(\mathcal{H}_{K})=H^{\infty}(\mathbb{D})$. If $\mathcal{E}_{T}\sim_{u}\mathcal{E}_{M_{z}^{*}}\otimes\mathcal{E}$ for some Hermitian holomorphic vector bundle $\mathcal{E}$ over $\mathbb{D}$ of rank $m$, then the following statements are equivalent:
@@
    There exist constants $C_{1}$ and $C_{2}$ such that the metric $h_{\mathcal{E}}(w)$ of $\mathcal{E}$ in some holomorphic frame satisfies $0<C_{1}I_{m}\leq h_{\mathcal{E}}(w)\leq C_{2}I_{m}$ for any $w\in\mathbb{D}.$
@@
    There exists a holomorphic frame $\gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ of $\mathcal{E}$ such that each $\|\gamma_{i}(w)\|,1\leq i\leq m$, is uniformly bounded on $\mathbb{D}$, and there are constants $A_{1}$ and $A_{2}$ such that the metric $h_{\mathcal{E}}(w)$ under the holomorphic frame $\gamma$ satisfies $0<A_{1}\leq\det h_{\mathcal{E}}(w)\leq A_{2}$ for any $w\in\mathbb{D}.$
In Proposition 3.6, if the metric of the holomorphic complex bundle $\mathcal{E}$ under a holomorphic frame is a polynomial for $|w|^{2}$, with each coefficient being a non-negative real number, then we can construct many contractive operators similar to the backward shift operators on the Hardy space or the weighted Bergman spaces. This provides positive support for Theorem 1 in [11].
@@ 4. Similarity on weighted Dirichlet spaces
Shields, in [49], provided a necessary and sufficient condition for the similarity of backward shift operators on Hilbert space using weight sequences. From [28, 40], we know when a tensor product of Hermitian holomorphic vector bundles corresponds to a bounded linear operator. The goal of this section is to use the holomorphic jet bundle to describe the similarity of the backward shift operator on weighted Dirichlet spaces.
Let $\mathcal{E}$ be a Hermitian holomorphic line bundle over $\mathbb{D}$ and $\gamma$ be a non-vanishing holomorphic cross-section of $\mathcal{E}$. For $k=0,1,2,\cdots$, we associate to $\mathcal{E}$ a $(k+1)$-dimensional Hermitian holomorphic vector bundle $\mathcal{J}_{k}(\mathcal{E})$ called the holomorphic $k$-jet bundle of $\mathcal{E}$. The work [13, 22, 35] are good references for $k$-jet bundles.
Suppose that for each $w$ in $\mathbb{D}$, $\gamma(w),\gamma^{{}^{\prime}}(w),\gamma^{(2)}(w),\ldots,\gamma^{(k)}(w)$ are independent. The $k$-jet bundle $\mathcal{J}_{k}(\mathcal{E})$ of the bundle $\mathcal{E}$ has an associated holomorphic frame $\mathcal{J}_{k}(\gamma)=\{\gamma,\gamma^{{}^{\prime}},\gamma^{(2)},\ldots,\gamma^{(k)}\}$. The metric $h_{\mathcal{E}}(w)=\langle\gamma(w),\gamma(w)\rangle$ for $\mathcal{E}$ induces a metric $\mathcal{J}_{k}(h_{\mathcal{E}})(w)$ for $\mathcal{J}_{k}(\mathcal{E})$ of the form
@@
Throughout the section, we denote by $D_{\alpha}^{*}$ and $D^{*}_{\alpha,E}$ the backward shift operators on weighted Dirichlet space $\mathcal{D}_{\alpha}$ and the space $\mathcal{D}_{\alpha,E}$, respectively, where $\alpha\in(0,1]$ and $E$ is a separable Hilbert space. The reproducing kernel of $\mathcal{D}_{\alpha}$ is $K_{\alpha}(z,w)=\sum\limits_{i=0}^{\infty}\frac{z^{i}\overline{w}^{i}}{(i+1)^{\alpha}},z,w\in\mathbb{D}$. In particular, $D_{1}^{*}$ and $\mathcal{D}_{1}$ are also denoted as $D^{*}$ and $\mathcal{D}$, respectively. Following the notations and ideas of Müller in [43], the Theorem 2.6, we use the following lemma to demonstrate how to construct a defect operator corresponding to operator $T\in\mathcal{L}(\mathcal{H})$ with $\sum\limits_{s=1}^{\infty}\frac{\|T^{s}\|^{2}}{(s+1)^{\alpha}}\leq 1$ such that an operator $T$ is unitary equivalent to the restriction of $D^{*}_{\alpha,E}$ on some invariant subspace.
@@
Let $b_{0}=1$, $c_{0}=1$, and for $k\geq 1,$ $b_{k}=\frac{1}{(k+1)^{\alpha}}$, and $c_{k}=-\sum\limits_{j=0}^{k-1}\frac{c_{j}}{(k-j+1)^{\alpha}}$, then for any fixed positive integer $l\geq 1$, $\sum\limits_{i=0}^{l}b_{i}c_{l-i}=0$. Obviously, $-b_{1}=c_{1}<0$. Without losing generality, assume that $-b_{i}\leq c_{i}<0$ for $11\leq i\leq k-$ 1. Then by the assumption $\sum\limits_{s=1}^{\infty}\frac{\|T^{s}\|^{2}}{(s+1)^{\alpha}}\leq 1$ and Lemma 2.1 in [43], we know that $0\geq c_{k}\geq-b_{k}$ for every $k\geq 1,$ which gives
@@
$\small\|Vx\|^{2}=\sum\limits_{n=0}^{\infty}b_{n}\left\langle\left(\sum\limits_{k=0}^{\infty}c_{k}(T^{k})^{*}T^{k}\right)T^{n}x,T^{n}x\right\rangle=\sum\limits_{l=0}^{\infty}\left(\|T^{l}x\|^{2}\sum\limits_{n+k=l}b_{n}c_{k}\right)=\|x\|^{2}$
@@
Hence, $T\sim_{u}D^{*}_{\alpha,E}|_{\mathcal{N}}$ for the $D_{\alpha,E}^{*}$-invariant subspace $\mathcal{N}\subset\mathcal{D}_{\alpha,E}$. ∎
Similar to the definition of the defect operator of $n$-hypercontraction in Section 2, to prevent confusion, we still denote $D_{T}=\left(\sum\limits_{k=0}^{\infty}c_{k}(T^{k})^{*}T^{k}\right)^{1/2}$ as the defect operator of the operator $T$ with $\sum\limits_{s=1}^{\infty}\frac{\|T^{s}\|^{2}}{(s+1)^{\alpha}}\leq 1$. Then we have the following lemma from Lemma 4.1, whose proof is omitted.
@@
The ratio of the metrics on holomorphic line bundles, as derived in the previous section, serves as a similarity invariant for the backward shift operators on the Hardy space and the weighted Bergman space. However, whether this extends to weighted Dirichlet remains unclear. In what follows, we give a sufficient condition for certain bounded linear operators to be similar to backward shift operator $D_{\alpha}^{*}$ on weighted Dirichlet space $\mathcal{D}_{\alpha}$, derived through the holomorphic jet bundle.
@@
Let $T\in\mathcal{B}_{1}(\mathbb{D})$ and $\sum\limits_{s=1}^{\infty}\frac{\|T^{s}\|^{2}}{(s+1)^{\alpha}}\leq 1$ for some constant $\alpha\in(0,1]$. If there exist constants $c_{2}>c_{1}>0$ satisfy
@@
Since $T\in\mathcal{B}_{1}(\mathbb{D})$ satisfies $\sum\limits_{s=1}^{\infty}\frac{\|T^{s}\|^{2}}{(s+1)^{\alpha}}\leq 1$ for some constant $\alpha\in(0,1]$, by Lemma 4.1 and Lemma 4.2, we obtain that $\mathcal{E}_{T}\sim_{u}\mathcal{E}_{D_{\alpha}^{*}}\otimes\mathcal{E}_{D_{T}}$ for the holomorphic line bundle $\mathcal{E}_{D_{T}}$ induced by the defect operator $D_{T}$.
@@
Since $\|zf\|_{\mathcal{D}_{\alpha}}^{2}=\sum\limits_{i=0}^{\infty}(i+2)^{\alpha}|a_{i}|^{2}\leq 2\sum\limits_{i=0}^{\infty}(i+1)^{\alpha}|a_{i}|^{2}<\infty$, $z\in\text{Mult}(\mathcal{D}_{\alpha})$. From [41], $\phi\in H^{\infty}(\mathbb{D})$ is in $\text{Mult}(\mathcal{D}_{\alpha})$ if and only if there is a constant $d>0$ such that
@@
for some constants $c_{2}>c_{1}>0$, where $h_{\mathcal{E}_{D_{T}}}$ denotes the metric obtained from a non-vanishing holomorphic cross-section $\gamma$ of $\mathcal{E}_{D_{T}}$. Furthermore, since $\mathcal{E}_{D_{T}}(w)\subset\mathbb{C}^{n}$ for some $n>0$, there is an orthonormal basis $\{e_{i}\}_{i=1}^{n}$ of the Hilbert space $E=\bigvee\limits_{w\in\mathbb{D}}\gamma(w)$ such that $\gamma(w)=\sum\limits_{i=1}^{n}\gamma_{i}(w)e_{i}$. From (4.2),
(4.3) $c_{1}\leq h_{\mathcal{E}_{D_{T}}}(w)=\sum\limits_{i=1}^{n}|\gamma_{i}(w)|^{2}<\sum\limits_{i=1}^{n}|\gamma_{i}(w)|^{2}+\sum\limits_{i=1}^{n}|\gamma^{{}^{\prime}}_{i}(w)|^{2}=h_{\mathcal{E}_{D_{T}}}(w)+\frac{\partial^{2}}{\partial w\partial\overline{w}}h_{\mathcal{E}_{D_{T}}}(w)\leq 2c_{2},\quad w\in\mathbb{D}.$
By (4.1) and (4.3), we have for every $f\in\mathcal{D}_{\alpha}$,
@@
Then for every $f\in\mathcal{D}_{\alpha}$, $M_{F}f=f\gamma$ and
@@
where $h=\sum\limits_{i=1}^{n}h_{i}e_{i}\in E.$ It is easy to see that $G\in\text{Mult}(\mathcal{D}_{\alpha,E},\mathcal{D}_{\alpha})$ and $(G^{\#})^{*}\in\text{Mult}(\mathcal{D}_{\alpha},\mathcal{D}_{\alpha,E})$ for $G^{\#}(w)=G(\overline{w})$. Hence, by Theorem 2.1, for the reproducing kernel $K_{\alpha}(z,w)=\sum\limits_{n=0}^{\infty}\frac{z^{n}\overline{w}^{n}}{(n+1)^{\alpha}}$ of $\mathcal{D}_{\alpha}$,

### v3

@@ 3. Similarity in the class $\mathcal{B}_{m}(\mathbb{D})$
where $E=\bigvee\limits_{w\in\mathbb{D}}\{\widetilde{\gamma}_{i}(w):1\leq i\leq m\}$. Obviously, the function $F^{\#}$ defined on $\mathbb{D}$ as $F^{\#}(w):=F(\overline{w})$ is such that $(F^{\#})^{*}\in\text{Mult}(\mathcal{M}_{n,E},\mathcal{M}_{n,\mathbb{C}^{m}}).$ Moreover, by Theorem 2.1,
@@
Theorem 2.4 in [16] and Theorem 2.3 in [24] use the curvature of holomorphic complex bundles to characterize the similarity of $n$-hypercontractions. Using Theorem 3.4, we are able to give the following quick, alternative proof:
@@
Let $T\in\mathcal{B}_{m}(\mathbb{D})$ be an $n$-hypercontraction. Then $T$ is similar to the backward shift operator $S^{*}_{n,\mathbb{C}^{m}}$ on $\mathcal{M}_{n,\mathbb{C}^{m}}$ if and only if there exists a bounded subharmonic function $\varphi$ over $\mathbb{D}$ such that
$\text{trace}\;\mathcal{K}_{S^{*}_{n,\mathbb{C}^{m}}}(w)-\text{trace}\;\mathcal{K}_{T}(w)=\frac{\partial^{2}}{\partial w\partial\overline{w}}\varphi(w),\quad w\in\mathbb{D}.$
@@
Since $T\in\mathcal{B}_{m}(\mathbb{D})$ is an $n$-hypercontraction, by Lemma 3.2, there exists a Hermitian holomorphic vector bundle $\mathcal{E}_{D_{T}}$ over $\mathbb{D}$ of rank $m$ such that $\mathcal{E}_{T}\sim_{u}\mathcal{E}_{S^{*}_{n}}\otimes\mathcal{E}_{D_{T}}$.
If $T$ is similar to $S^{*}_{n,\mathbb{C}^{m}}$, by Theorem 3.4, there exist constants $A_{1}$, $A_{2}$, and a holomorphic frame $\gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ of $\mathcal{E}_{D_{T}}$ such that the metric $h_{\mathcal{E}_{D_{T}}}$ under $\gamma$ satisfies
$0<A_{1}\leq\det h_{\mathcal{E}_{D_{T}}}(w)\leq A_{2},\quad w\in\mathbb{D}.$
Setting $\det\mathcal{E}_{D_{T}}$ as the determinant bundle of $\mathcal{E}_{D_{T}}$, the metric on the determinant bundle is $h_{\det\mathcal{E}_{D_{T}}}=\det h_{\mathcal{E}_{D_{T}}},$ and we have
$\text{trace}\;\mathcal{K}_{S^{*}_{n,\mathbb{C}^{m}}}(w)-\text{trace}\;\mathcal{K}_{T}(w)=-\text{trace}\;\mathcal{K}_{\mathcal{E}_{D_{T}}}(w)=-\mathcal{K}_{\det\mathcal{E}_{D_{T}}}(w)=\frac{\partial^{2}}{\partial w\partial\overline{w}}\log h_{\det\mathcal{E}_{D_{T}}}(w)\geq 0.$
To complete the proof, one sets $\varphi(w):=\log h_{\text{det}\mathcal{E}_{D_{T}}}(w)$ on $\mathbb{D}$.
For the other implication, note that since $\mathcal{E}_{T}\sim_{u}\mathcal{E}_{S^{*}_{n}}\otimes\mathcal{E}_{D_{T}}$, we have
$\text{trace}\;\mathcal{K}_{T}=\text{trace}\;\mathcal{K}_{S^{*}_{n,\mathbb{C}^{m}}}+\text{trace}\;\mathcal{K}_{\mathcal{E}_{D_{T}}}.$
Thus,
$\frac{\partial^{2}}{\partial w\partial\overline{w}}\varphi(w)=-\text{trace}\;\mathcal{K}_{\mathcal{E}_{D_{T}}}(w)=-\mathcal{K}_{\det\mathcal{E}_{D_{T}}}(w)=\frac{\partial^{2}}{\partial w\partial\overline{w}}\log\|{\gamma}(w)\|^{2}$
for a bounded subharmonic function $\varphi$ on $\mathbb{D}$ and a non-vanishing holomorphic frame ${\gamma}$ of the determinant bundle $\det\mathcal{E}_{D_{T}}$. Now, $u(w):=\log\left(\frac{\|{\gamma}(w)\|^{2}}{e^{\varphi(w)}}\right)^{\frac{1}{2}}$ is a real-valued harmonic function, and letting $\phi:=e^{u+iv}\in\text{Hol}(\mathbb{D}),$ where $v$ is the harmonic conjugate of $u$, we see that
@@
Hence, $e^{{\varphi}(w)}=\left\|\frac{1}{\phi(w)}{\gamma}(w)\right\|^{2}$ and $\frac{1}{\phi}{\gamma}$ is a holomorphic frame of $\det\mathcal{E}_{D_{T}}$. Since $\varphi$ is bounded on $\mathbb{D}$, there are constants $C_{1}$ and $C_{2}$ such that
@@
and therefore by Theorem 3.4 and Theorem 3.2 in [38], $T\sim_{s}S^{*}_{n,\mathbb{C}^{m}}$. ∎
Two Cowen-Douglas operators are unitarily equivalent if and only if their corresponding holomorphic complex bundles are equivalent as Hermitian holomorphic vector bundles. It is known from [40] that many holomorphic complex bundles with tensor products correspond to some Cowen-Douglas operator. Therefore, we use holomorphic complex bundles to give a similarity classification of Cowen-Douglas operators and obtain the following proposition. We omit the proof of it which is similar to those of Theorem 3.4 and Corollary 3.5:
@@
Let $T\in\mathcal{B}_{m}(\mathbb{D})$ and let $M_{z}^{*}\in\mathcal{B}_{1}(\mathbb{D})$ be the backward shift operator on a reproducing kernel Hilbert space $\mathcal{H}_{K}$ with $\text{Mult}(\mathcal{H}_{K})=H^{\infty}(\mathbb{D})$. If $\mathcal{E}_{T}\sim_{u}\mathcal{E}_{M_{z}^{*}}\otimes\mathcal{E}$ for some Hermitian holomorphic vector bundle $\mathcal{E}$ over $\mathbb{D}$ of rank $m$, then the following statements are equivalent:
@@
    There exist constants $C_{1}$ and $C_{2}$ such that the metric $h_{\mathcal{E}}(w)$ of $\mathcal{E}$ in some holomorphic frame satisfies $0<C_{1}I_{m}\leq h_{\mathcal{E}}(w)\leq C_{2}I_{m}$ for every $w\in\mathbb{D}.$
@@
    There exists a holomorphic frame $\gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ of $\mathcal{E}$ such that each $\|\gamma_{i}(w)\|,1\leq i\leq m$, is uniformly bounded on $\mathbb{D}$, and there are constants $A_{1}$ and $A_{2}$ such that the metric $h_{\mathcal{E}}(w)$ under the holomorphic frame $\gamma$ satisfies
    $0<A_{1}\leq\det h_{\mathcal{E}}(w)\leq A_{2},\quad w\in\mathbb{D}.$
    In other words, the metric $h_{\det\mathcal{E}}$ of the determinant bundle $\det\mathcal{E}$ satisfies $0<A_{1}\leq h_{\det\mathcal{E}}(w)\leq A_{2}$ for every $w\in\mathbb{D}.$
  • (4)
    There exists a bounded subharmonic function $\varphi$ over $\mathbb{D}$ such that
    $m\mathcal{K}_{M_{z}^{*}}(w)-\text{trace}\;\mathcal{K}_{T}(w)=\frac{\partial^{2}}{\partial w\partial\overline{w}}\varphi(w),\quad w\in\mathbb{D}.$
In Proposition 3.6, if the metric of the holomorphic complex bundle $\mathcal{E}$ under a holomorphic frame is a polynomial for $|w|^{2}$, with each coefficient being a non-negative real number, then we can construct numerous examples of contractive operators similar to the backward shift operators on the Hardy or weighted Bergman spaces. These examples satisfy the conditions of Theorem 1 in [11].
@@ 4. Similarity on weighted Dirichlet spaces
Shields, in [49], provided a necessary and sufficient condition for the similarity of backward shift operators on Hilbert space using weight sequences. From [28, 40], we know when a tensor product of Hermitian holomorphic vector bundles corresponds to a bounded linear operator. The goal of this section is to use the holomorphic jet bundle to describe the similarity of backward shifts on weighted Dirichlet spaces.
Let $\mathcal{E}$ be a Hermitian holomorphic line bundle over $\mathbb{D}$ and let $\gamma$ be a non-vanishing holomorphic cross-section of $\mathcal{E}$. For $k=0,1,2,\cdots$, we associate to $\mathcal{E}$ a $(k+1)$-dimensional Hermitian holomorphic vector bundle $\mathcal{J}_{k}(\mathcal{E})$ called the holomorphic $k$-jet bundle of $\mathcal{E}$. The work [13, 22, 35] are good references for $k$-jet bundles. Suppose that for each $w\in\mathbb{D}$, $\gamma(w),\gamma^{{}^{\prime}}(w),\gamma^{(2)}(w),\ldots,\gamma^{(k)}(w)$ are independent. The $k$-jet bundle $\mathcal{J}_{k}(\mathcal{E})$ of the bundle $\mathcal{E}$ has an associated holomorphic frame $\mathcal{J}_{k}(\gamma)=\{\gamma,\gamma^{{}^{\prime}},\gamma^{(2)},\ldots,\gamma^{(k)}\}$. The metric $h_{\mathcal{E}}(w)=\langle\gamma(w),\gamma(w)\rangle$ for $\mathcal{E}$ induces a metric $\mathcal{J}_{k}(h_{\mathcal{E}})(w)$ for $\mathcal{J}_{k}(\mathcal{E})$ of the form
@@
Throughout the section, we denote by $D_{\alpha}^{*}$ and $D^{*}_{\alpha,E}$ the backward shift operators on weighted Dirichlet spaces $\mathcal{D}_{\alpha}$ and the spaces $\mathcal{D}_{\alpha,E}$, respectively, where $\alpha\in(0,1]$ and $E$ is a separable Hilbert space. The reproducing kernel of $\mathcal{D}_{\alpha}$ is given by $K_{\alpha}(z,w)=\sum\limits_{i=0}^{\infty}\frac{z^{i}\overline{w}^{i}}{(i+1)^{\alpha}},z,w\in\mathbb{D}$. In particular, $D_{1}^{*}$ is also denoted as $D^{*}$ and $\mathcal{D}_{1}$ as $\mathcal{D}$.
We will use the following result by Müller in [43], Theorem 2.6, to construct a defect operator corresponding to $T\in\mathcal{L}(\mathcal{H})$ with $\sum\limits_{s=1}^{\infty}\frac{\|T^{s}\|^{2}}{(s+1)^{\alpha}}\leq 1$ such that $T$ is unitary equivalent to the restriction of $D^{*}_{\alpha,E}$ on some invariant subspace:
@@
Let $b_{0}=1$, $c_{0}=1$, and for $k\geq 1,$ set $b_{k}=\frac{1}{(k+1)^{\alpha}}$, and $c_{k}=-\sum\limits_{j=0}^{k-1}\frac{c_{j}}{(k-j+1)^{\alpha}}$. Then for a fixed positive integer $l\geq 1$, $\sum\limits_{i=0}^{l}b_{i}c_{l-i}=0$. Obviously, $-b_{1}=c_{1}<0$. Without losing generality, assume that $-b_{i}\leq c_{i}<0$ for $11\leq i\leq k-$ 1. Then by the assumption $\sum\limits_{s=1}^{\infty}\frac{\|T^{s}\|^{2}}{(s+1)^{\alpha}}\leq 1$ and Lemma 2.1 in [43], we know that $0\geq c_{k}\geq-b_{k}$ for every $k\geq 1,$ which gives
@@
$\small\|Vx\|^{2}=\sum\limits_{n=0}^{\infty}b_{n}\left\langle\left(\sum\limits_{k=0}^{\infty}c_{k}(T^{k})^{*}T^{k}\right)T^{n}x,T^{n}x\right\rangle=\sum\limits_{l=0}^{\infty}\left(\|T^{l}x\|^{2}\sum\limits_{n+k=l}b_{n}c_{k}\right)=\|x\|^{2},$
@@
Hence, $T\sim_{u}D^{*}_{\alpha,E}|_{\mathcal{N}}$ where $\mathcal{N}\subset\mathcal{D}_{\alpha,E}$ is a $D_{\alpha,E}^{*}$-invariant subspace. ∎
We will denote as
$D_{T}=\left(\sum\limits_{k=0}^{\infty}c_{k}(T^{k})^{*}T^{k}\right)^{1/2},$
the defect operator of the operator $T$ with $\sum\limits_{s=1}^{\infty}\frac{\|T^{s}\|^{2}}{(s+1)^{\alpha}}\leq 1$. Then we have the following lemma from Lemma 4.1, whose proof is omitted.
@@
The ratio of the metrics on holomorphic line bundles, as derived in the previous section, serves as a similarity invariant for the backward shift operators on the Hardy and weighted Bergman spaces. However, whether this extends to weighted Dirichlet spaces remains unclear. In what follows, we give a sufficient condition for certain bounded linear operators to be similar to the backward shift operator $D_{\alpha}^{*}$ on weighted Dirichlet space $\mathcal{D}_{\alpha}$ in terms of the holomorphic jet bundle.
@@
Let $T\in\mathcal{B}_{1}(\mathbb{D})$ and $\sum\limits_{s=1}^{\infty}\frac{\|T^{s}\|^{2}}{(s+1)^{\alpha}}\leq 1$ for some constant $\alpha\in(0,1]$. If there exist constants $c_{2}>c_{1}>0$ that satisfy
@@
Since $\sum\limits_{s=1}^{\infty}\frac{\|T^{s}\|^{2}}{(s+1)^{\alpha}}\leq 1$ for some constant $\alpha\in(0,1]$, by Lemmas 4.1 and 4.2, $\mathcal{E}_{T}\sim_{u}\mathcal{E}_{D_{\alpha}^{*}}\otimes\mathcal{E}_{D_{T}}$ for the holomorphic line bundle $\mathcal{E}_{D_{T}}$ induced by the defect operator $D_{T}$.
@@
Since $\|zf\|_{\mathcal{D}_{\alpha}}^{2}=\sum\limits_{i=0}^{\infty}(i+2)^{\alpha}|a_{i}|^{2}\leq 2\sum\limits_{i=0}^{\infty}(i+1)^{\alpha}|a_{i}|^{2}<\infty$, $z\in\text{Mult}(\mathcal{D}_{\alpha})$. From [41], it is known that $\phi\in H^{\infty}(\mathbb{D})$ is in $\text{Mult}(\mathcal{D}_{\alpha})$ if and only if there is a constant $d>0$ such that
@@
for some constants $c_{2}>c_{1}>0$, where $h_{\mathcal{E}_{D_{T}}}$ denotes the metric obtained from a non-vanishing holomorphic cross-section $\gamma$ of $\mathcal{E}_{D_{T}}$. Furthermore, since $\mathcal{E}_{D_{T}}(w)\subset\mathbb{C}^{n}$ for some $n>0$, there is an orthonormal basis $\{e_{i}\}_{i=1}^{n}$ of the Hilbert space $E=\bigvee\limits_{w\in\mathbb{D}}\gamma(w)$ such that $\gamma(w)=\sum\limits_{i=1}^{n}\gamma_{i}(w)e_{i}$. From (4.2), it follows that
(4.3) $c_{1}\leq h_{\mathcal{E}_{D_{T}}}(w)=\sum\limits_{i=1}^{n}|\gamma_{i}(w)|^{2}<\sum\limits_{i=1}^{n}|\gamma_{i}(w)|^{2}+\sum\limits_{i=1}^{n}|\gamma^{{}^{\prime}}_{i}(w)|^{2}=h_{\mathcal{E}_{D_{T}}}(w)+\frac{\partial^{2}}{\partial w\partial\overline{w}}h_{\mathcal{E}_{D_{T}}}(w)\leq c_{2},\quad w\in\mathbb{D}.$
Then by (4.1) and (4.3), we have for every $f\in\mathcal{D}_{\alpha}$,
@@
Then for every $f\in\mathcal{D}_{\alpha}$, $M_{F}f=f\gamma$, and
@@
where $h=\sum\limits_{i=1}^{n}h_{i}e_{i}\in E.$ It is easy to see that $G\in\text{Mult}(\mathcal{D}_{\alpha,E},\mathcal{D}_{\alpha})$ and $(G^{\#})^{*}\in\text{Mult}(\mathcal{D}_{\alpha},\mathcal{D}_{\alpha,E})$ for $G^{\#}(w)=G(\overline{w})$. Hence, using Theorem 2.1 and the reproducing kernel $K_{\alpha}(z,w)=\sum\limits_{n=0}^{\infty}\frac{z^{n}\overline{w}^{n}}{(n+1)^{\alpha}}$ of $\mathcal{D}_{\alpha}$,